A 62.25gm sample of CuSO4.5H2O(MOl wt-249.7) is dissolved in enough water to make 0.80L of solution.What volume of this solution must be diluted with H2O to make 1L of o.1M CuSO4.


pls explain with steps. its urgent 

Grade:12

1 Answers

Sakshi
8 years ago
First of all, one should calculate the molarity of the solution so prepared.
Molarity of the solution 1 = 62.25/(249.7*0.8) = 0.311 M
On applying the dilution formula as M1* V1 = M2 * V2
0.311* V= 0.1 * 1L
One can calculate the value of V as the answer